If roots of two equations are in AP find ratio of their discriminants

Grade:11

1 Answers

BALAJI ANDALAMALA
8 years ago
Let a,b and c,d are the roots of the two equations.
The equatios are
x^2-(a+b)x+ab=0\,\,and\,\,x^2-(c+d)x+cd=0

With discriminants D1 and D2.

D_1 = (a+b)^2-4ab = (a-b)^2
D_2 = (c+d)^2-4cd = (c-d)^2
Given a,b,c,d are in A.P , So b-a = d-c

\therefore \frac{D_1}{D_2} = \frac{(a-b)^2}{(c-d)^2} = 1